# TransformerFAM: Feedback attention is working memory

### v2

@@ C.1 Beyond Block-Recurrent Transformers
Block-Recurrent Transformers (BRT) (Hutchins et al., 2022) have a recurrent state that connects each block, and the recurrent state and the block input sequence exchange information through cross-attention. In this architecture, the recurrent state plays a role similar to FAM. We started the project by removing the additional QKV projection and cross-attention for the recurrent state in BRT and integrating all computations into self-attention.
@@ C.2 Feedback Memory Segment
As shown in Figure 2(b), TransformerFAM utilizes the updated FAM from the previous block as the input to the current block. Therefore, it is natural to consider using the Transformer outputs of the previous block directly as the input to the current block, instead of using a complex FAM mechanism. Figure 9 illustrates this modification, which we refer to as the Feedback Memory Segment (FM). FM achieves an infinite theoretical receptive field by performing self-attention at the output level. This is a specific variant of Staircase Attention (Ju et al., 2022). Staircase Attention proposes using activations from progressively higher Transformer layers as the memory segment goes further into the past.
@@
However, as shown in Figure 10, FM fails to retain PassKey information over a long context. The figure compares M1, 2, and 4 with FM1, 2, and 4 when the memory segment size is 1, 2, and 4 for BSWA and Feedback Memory Segment. TransformerFM outperforms TransformerBSWA, but still falls short of TransformerFAM by a significant margin. In TransformerFM, each activation must possess both local and global representation, similar to TransformerBSWA. However, the absence of an activation specifically responsible for the global representation appears to prevent the retention of critical information like PassKey over an extended context.
@@ C.3 Static Summary Tokens
As illustrated in Figure 2(b), ”FAM copy” duplicates the previously updated FAM to the current FAM input. However, ”FAM copy” is not an essential component in the feedback loop. When FAM compresses the current block, it can also attend to the previous FAM as both key and value, as shown in the experiment in Section 3.4.2. Therefore, ”FAM copy” is not strictly necessary because past FAM information is propagated through self-attention.
As mentioned in Section B.1, we learn the initial FAM using prompt tuning, and this token can be considered a ”Summary” token. An alternative design could employ a persistent summary token for block compression, with the feedback loop enabled by attending to the key-value pairs of previous FAMs.
As illustrated in Figure 10, SUM FAM (static summary token) prevents the model from successfully solving the PassKey task. This suggests that the query for summarization needs to be dynamically generated conditioned on the past FAM. A static query may not be able to cover all situations effectively. Additionally, simply attending to the key and value alone does not transmit sufficient information from the past. On the other hand, the ”FAM copy” mechanism propagates information from the past FAM to the current FAM through the residual connection of the Transformer, which facilitates better performance on the PassKey task.
@@ C.4 Diversity Loss
In Section 3.4.1, we observed a performance drop when the FAM length exceeded 64. We hypothesized that FAM was underutilizing its capacity and experiencing mode collapse, where lengthy FAM either focused on a small number of inputs or prevented all inputs from obtaining sufficient information.
As a remedy, we employed diversity loss as an auxiliary loss, similar to Wav2vec 2.0 (Baevski et al., 2020). Diversity loss aims to maximize the entropy of QK attention probability, as shown in Equation 2b. This loss encourages FAM to uniformly attend to all inputs, and all inputs equally attend to FAM. In Equation 2, $b$ denotes the batch size, $\tau$ represents the block index, $l$ indicates the sequence position within a block, and $h$ refers to multi heads. In Equation 2a, $\bar{p}_{b\tau}$ represents the average attention probability across all sequences and multi-heads for the self-attention of each block. The auxiliary loss functions to regularize this probability towards uniformity.
@@ C.5 Reconstruction Loss
Compressive transformers (Rae et al., 2019) compress memory segments before forwarding them to the next block. The key difference between our approach is that the compressed information is not recurrently connected. The paper proposes a reconstruction loss as an auxiliary loss, which aims to reconstruct the original activations from the compressed activations.
@@ D.1 PassKey Retrieval
The format of the PassKey retrieval task is shown in Figure 11. The original paper (Mohtashami & Jaggi, 2023) also included a prefix filler, but we removed it in our paper.
@@
Figure 11: PassKey Retrieval Format (Mohtashami & Jaggi, 2023)
@@ D.3 GPT-3 Tasks
We evaluated all model sizes on the tasks reported by GPT-3 (Brown et al., 2020). The results are shown in Table 12.
@@
Table 12: Comparison of performance of GPT-3 tasks (Brown et al., 2020) between BSWA and FAM across 1B, 8B, and 24B models.
@@ Appendix E Related Work
The Transformer architecture exhibits a quadratic complexity with respect to context length, a significant limitation. To address this, several research works have focused on approximating the attention mechanism. One approach involves sparse attention, where only a subset of important tokens are attended to, as seen in models like Sparse Transformer (Child et al., 2019), Big Bird (Zaheer et al., 2020), Reformer (Kitaev et al., 2020), Routing Transformer (Roy et al., 2021), and TOVA (Oren et al., 2024). Linear approximation methods offer an alternative, seeking to circumvent the quadratic complexity by altering attention calculations, as exemplified by Linformer (Wang et al., 2020), Linear Transformer (Katharopoulos et al., 2020), Performer (Choromanski et al., 2020), and Nyströmformer (Xiong et al., 2021). Finally, some research explores entirely different sequence-to-sequence architectures as replacements for attention-based Transformers, including MLP-mixer (Tolstikhin et al., 2021), State Space Models (Rangapuram et al., 2018), S4 (Gu et al., 2021), Mamba (Gu & Dao, 2023), and RWKV (Peng et al., 2023).
Global Workspace Theory (GWT) (Baars, 2005) is a leading theory of consciousness. According to GWT, the human brain possesses a global workspace where various modules, such as sensory input, memories, and internal representations, converge. The attention mechanism in brain acts as a spotlight, focusing on specific inputs among the multitude, and transforming this unconscious activity into conscious awareness. These ”momentarily active, subjectively experienced” events are then stored in working memory. TransformerFAM draws inspiration from GWT, adopting its principle of a unified attention mechanism for processing homogenous, heterogeneous, and feedback data.
@@ Appendix F Attention Visualization
Figure 12 depicts the attention map for each head in each layer of a 1B model. FAM is prepended and is located at the bottom left corner. The bright spots along the left edge represent the block inputs attending to FAM, while the bright spots along the bottom edge represent FAM compressing the corresponding block. Overall, the block inputs actively reference FAM, while FAM compresses only the selective inputs.

### v3

@@ C.1 Beyond Block-Recurrent Transformers
Block-Recurrent Transformers (BRT) [52] have a recurrent state that connects each block, and the recurrent state and the block input sequence exchange information through cross-attention. In this architecture, the recurrent state plays a role similar to FAM. We started the project by removing the additional QKV projection and cross-attention for the recurrent state in BRT and integrating all computations into self-attention.
@@
Around the time this paper was published, Infini-Transformer [79] was also released. Infini-Transformer reported successful PassKey retrieval despite the compressed information not passing through the FF layer. This is because it stores weighted values in scratchpad memory and reuses them, preserving the latent space of the values. However, since the weighted values are all it can represent, there might be limitations in its expressiveness.
Furthermore, TransformerFAM maintains the past memory segment from BSWA despite having compressed memory because the compressed memory cannot retain detailed information. Since Infini-Transformer completely discards the past memory segment, it might have difficulty remembering recent details.
@@ C.2 Feedback Memory Segment
As shown in Fig. 2(b), TransformerFAM utilizes the updated FAM from the previous block as the input to the current block. Therefore, it is natural to consider using the Transformer outputs of the previous block directly as the input to the current block, instead of using a complex FAM mechanism. Fig. 8 illustrates this modification, which we refer to as the Feedback Memory Segment (FM). FM achieves an infinite theoretical receptive field by performing self-attention at the output level. This is similar to ERNIE-Doc [30], and a specific variant of Staircase Attention [61]. Staircase Attention proposes using activations from progressively higher Transformer layers as the memory segment goes further into the past.
@@
However, as shown in Fig. 9, FM fails to retain PassKey information over a long context. The figure compares M1, 2, and 4 with FM1, 2, and 4 when the memory segment size is 1, 2, and 4 for BSWA and Feedback Memory Segment. TransformerFM outperforms TransformerBSWA, but still falls short of TransformerFAM by a significant margin. In TransformerFM, each activation must possess both local and global representation, similar to TransformerBSWA. However, the absence of an activation specifically responsible for the global representation appears to prevent the retention of critical information like PassKey over an extended context.
@@ C.3 Static Summary Tokens
As illustrated in Fig. 2(b), "FAM copy" duplicates the previously updated FAM to the current FAM input. However, "FAM copy" is not an essential component in the feedback loop. When FAM compresses the current block, it can also attend to the previous FAM as both key and value, as shown in the experiment in Section B.4.2. Therefore, "FAM copy" is not strictly necessary because past FAM information is propagated through self-attention.
As mentioned in Section B.1, we learn the initial FAM using prompt tuning, and this token can be considered a "Summary" token. An alternative design could employ a persistent summary token for block compression, with the feedback loop enabled by attending to the key-value pairs of previous FAMs.
As illustrated in Fig. 9, SUM FAM (static summary token) prevents the model from successfully solving the PassKey task. This suggests that the query for summarization needs to be dynamically generated conditioned on the past FAM. A static query may not be able to cover all situations effectively. Additionally, simply attending to the key and value alone does not transmit sufficient information from the past. On the other hand, the "FAM copy" mechanism propagates information from the past FAM to the current FAM through the residual connection of the Transformer, which facilitates better performance on the PassKey task.
@@ C.4 Diversity Loss
In Section B.4.1, we observed a performance drop when the FAM length exceeded 64. We hypothesized that FAM was underutilizing its capacity and experiencing mode collapse, where lengthy FAM either focused on a small number of inputs or prevented all inputs from obtaining sufficient information.
As a remedy, we employed diversity loss as an auxiliary loss, similar to Wav2vec 2.0 [41]. Diversity loss aims to maximize the entropy of QK attention probability, as shown in Eq. 2b. This loss encourages FAM to uniformly attend to all inputs, and all inputs equally attend to FAM. In Eq. 2, $b$ denotes the batch size, $\tau$ represents the block index, $l$ indicates the sequence position within a block, and $h$ refers to multi heads. In Eq. 2a, $\bar{p}_{b\tau}$ represents the average attention probability across all sequences and multi-heads for the self-attention of each block. The auxiliary loss functions to regularize this probability towards uniformity.
@@ C.5 Reconstruction Loss
Compressive transformers [21] compress memory segments before forwarding them to the next block. The key difference between our approach is that the compressed information is not recurrently connected. The paper proposes a reconstruction loss as an auxiliary loss, which aims to reconstruct the original activations from the compressed activations.
@@ D.1 PassKey Retrieval
The format of the PassKey retrieval task is shown in Fig. 10. The original paper [71] also included a prefix filler, but we removed it in our paper.
@@
Figure 10: PassKey Retrieval Format [71]
@@ D.3 GPT-3 Tasks
We evaluated all model sizes on the tasks reported by GPT-3 [33]. The results are shown in Table 12.
@@
Table 12: Comparison of performance of GPT-3 tasks [33] between BSWA and FAM across 1B, 8B, and 24B models.
@@ Appendix E Related Work
The Transformer architecture exhibits a quadratic complexity with respect to context length, a significant limitation. To address this, several research works have focused on approximating the attention mechanism. One approach involves sparse attention, where only a subset of important tokens are attended to, as seen in models like Sparse Transformer [23], Big Bird [28], Reformer [37], Routing Transformer [42], and TOVA [81]. Linear approximation methods offer an alternative, seeking to circumvent the quadratic complexity by altering attention calculations, as exemplified by Linformer [34], Linear Transformer [40], Performer [38], and Nyströmformer [48]. Finally, some research explores entirely different sequence-to-sequence architectures as replacements for attention-based Transformers, including MLP-mixer [47], State Space Models [18], S4 [43], Mamba [73], and RWKV [76].
Global Workspace Theory (GWT) [8] is a leading theory of consciousness. According to GWT, the human brain possesses a global workspace where various modules, such as sensory input, memories, and internal representations, converge. The attention mechanism in brain acts as a spotlight, focusing on specific inputs among the multitude, and transforming this unconscious activity into conscious awareness. These "momentarily active, subjectively experienced" events are then stored in working memory. TransformerFAM draws inspiration from GWT, adopting its principle of a unified attention mechanism for processing homogenous, heterogeneous, and feedback data.
@@ Appendix F Attention Visualization
Fig. 11 depicts the attention map for each head in each layer of a 1B model. FAM is prepended and is located at the bottom left corner. The bright spots along the left edge represent the block inputs attending to FAM, while the bright spots along the bottom edge represent FAM compressing the corresponding block. Overall, the block inputs actively reference FAM, while FAM compresses only the selective inputs.
@@
Appendix G Limitations
While the results presented in Table 1 demonstrate that TransformerFAM shows improvements on long-context tasks, these gains are not yet substantial, highlighting the need for further development and refinement of working memory mechanisms.
In this paper, we have taken an initial step towards integrating attention-based working memory, a concept inspired by neuroscience, into deep learning architectures. We believe there is significant potential for further exploration in this direction, and we encourage future research to continue addressing the ongoing challenge of limited memory in deep learning models.
Appendix H Broader Impacts
While our work is inspired by the concept of working memory from neuroscience, as discussed in Section Section 3.3, achieving a human-level implementation remains a significant challenge. This research represents an initial step in that direction.
The potential societal impacts of such advanced working memory in LLMs could be substantial, with applications like highly personalized AI assistants. However, these impacts are currently speculative due to the nascent stage of this research.
In the immediate future, the primary benefit of our work is expected to be improvements in the efficiency and effectiveness of LLMs, with potential applications across various domains such as education, healthcare, and communication. As with any technology, we recognize the possibility of misuse and encourage ongoing research into the ethical implications of LLMs and related advancements in artificial intelligence.